# Leveraging Distribution Matching to Make Approximate Machine Unlearning Faster

### v2

@@ 4.2 The Efficiency and Efficacy of Blend Condensation
This dramatic speed-up does not compromise performance. As shown in Figure 2, Blend consistently yields higher test accuracy (e.g., 35.03% at IPC=10) compared to GM (33.01%), DM (32.2%), and IDM (33.1%). Furthermore, the synthetic prototypes generated by Blend are qualitatively superior—they are noticeably smoother and more class-consistent (e.g., clear airplane outlines) than the noisy, cluttered images from other methods. This demonstrates that our approach of grouping and blending feature-similar images is highly effective at producing coherent and informative exemplars.
@@
Random Class Forgetting (Condensation Time for 1000 IPC: $47.7812\pm 1.3323$ s)
Retraining $49.76\pm 2.57$ $50.51\pm 3.55$ $100.00\pm 0.00$ $75.50\pm 5.63$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $80.35\pm 2.08$ $62.47\pm 4.88$ $1141.71\pm 51.30$ $224.50\pm 0.14$
CF $50.18\pm 1.16$ $55.39\pm 2.67$ $100.00\pm 0.00$ $96.39\pm 0.75$ $00.00\pm 0$ 0.00 $0.30\pm 0.45$ $81.19\pm 2.40$ $77.75\pm 1.25$ $577.81\pm 55.36$ $112.32\pm 0.04$
@@
$L_{1}$ Sparse $47.32\pm 2.73$ $53.79\pm 0.88$ $99.00\pm 0.97$ $96.36\pm 0.55$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $78.42\pm 3.40$ $77.66\pm 1.32$ $358.62\pm 28.80$ $69.88\pm 0.01$
SCRUB $50.12\pm 0.37$ $50.00\pm 0.00$ $15.28\pm 7.66$ $11.11\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $13.80\pm 6.90$ $10.00\pm 0.00$ $241.23\pm 23.22$ $61.58\pm 0.02$
Pruning $50.87\pm 3.22$ $50.80\pm 2.27$ $98.08\pm 0.73$ $100.00\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $77.43\pm 2.41$ $72.65\pm 1.27$ $235.50\pm 6.20$ $46.74\pm 0.01$
A-CF $51.71\pm 2.29$ $48.25\pm 7.31$ $97.88\pm 4.39$ $98.92\pm 0.62$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $77.01\pm 3.20$ $77.45\pm 3.63$ $75.76\pm 0.27$ $16.01\pm 0.96$
10% Uniform Forgetting (Condensation Time for 1000 IPC: $34.1780\pm 0.7403$ s)
Retraining $50.47\pm 0.61$ $50.05\pm 1.64$ $100.00\pm 0.00$ $91.48\pm 0.41$ $87.48\pm 1.01$ $82.25\pm 1.01$ $86.84\pm 0.62$ $80.65\pm 0.48$ $1132.22\pm 33.71$ $664.73\pm 6.16$
@@ 4.3 Single-Round Unlearning
[…]
@@
Random Class Forgetting (Condensation Time for 500 IPC: $4.90615\pm 0.1941$ s)
Retraining $50.45\pm 1.53$ $52.01\pm 5.38$ $99.99\pm 0.00$ $78.46\pm 3.28$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $82.71\pm 10.54$ $68.53\pm 8.79$ $1012.11\pm 4.18$ $22.79\pm 0.04$
CF $50.34\pm 3.97$ $51.15\pm 0.32$ $100.00\pm 0.00$ $91.37\pm 0.40$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $82.37\pm 10.30$ $81.30\pm 0.25$ $505.95\pm 2.05$ $11.36\pm 0.01$
Bad Distillation $55.38\pm 1.74$ $49.58\pm 6.42$ $98.49\pm 2.02$ $99.69\pm 0.72$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $81.84\pm 9.47$ $83.04\pm 9.03$ $248.95\pm 1.34$ $45.15\pm 0.03$
$L_{1}$ Sparse $53.71\pm 3.01$ $52.07\pm 1.39$ $99.16\pm 0.96$ $96.97\pm 0.13$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $84.04\pm 1.87$ $94.18\pm 0.19$ $314.88\pm 3.28$ $7.14\pm 0.00$
SCRUB $50.47\pm 1.64$ $49.97\pm 0.03$ $12.93\pm 2.81$ $11.11\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $13.61\pm 8.37$ $7.37\pm 0.71$ $211.82\pm 2.00$ $19.57\pm 0.02$
Pruning $49.77\pm 2.35$ $51.90\pm 2.20$ $99.04\pm 0.18$ $100.00\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $80.80\pm 10.02$ $77.36\pm 8.78$ $210.93\pm 1.00$ $4.80\pm 0.01$
A-CF $52.04\pm 2.70$ $54.09\pm 6.30$ $99.65\pm 0.59$ $97.55\pm 0.71$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $80.90\pm 9.38$ $84.49\pm 2.80$ $83.27\pm 0.48$ $2.07\pm 0.02$
10% Uniform Forgetting (Condensation Time for 100 IPC: $0.4733\pm 0.0759$ s)
Retraining $51.75\pm 1.91$ $51.61\pm 3.54$ $99.99\pm 0.00$ $99.76\pm 0.03$ $94.60\pm 0.59$ $94.71\pm 0.26$ $95.18\pm 0.16$ $95.12\pm 0.17$ $1011.96\pm 6.45$ $992.98\pm 7.97$
@@
Bad Distillation $92.23\pm 0.43$ $93.57\pm 0.44$ $96.94\pm 1.02$ $99.75\pm 0.11$ $23.77\pm 8.50$ $23.77\pm 2.00$ $89.73\pm 2.44$ $91.72\pm 0.41$ $248.75\pm 0.95$ $369.65\pm 2.02$
$L_{1}$ Sparse $53.55\pm 1.96$ $53.12\pm 2.44$ $99.32\pm 0.44$ $99.47\pm 0.21$ $93.80\pm 1.00$ $94.32\pm 0.87$ $93.60\pm 0.57$ $93.33\pm 0.20$ $314.80\pm 2.90$ $308.10\pm 1.58$
SCRUB $54.99\pm 1.47$ $56.51\pm 4.66$ $99.66\pm 0.07$ $99.99\pm 0.02$ $95.00\pm 0.29$ $99.07\pm 0.44$ $94.48\pm 0.31$ $95.13\pm 0.19$ $211.64\pm 0.85$ $214.29\pm 1.76$
Pruning $54.20\pm 1.60$ $53.55\pm 0.72$ $99.33\pm 0.10$ $99.18\pm 0.40$ $94.91\pm 0.16$ $94.94\pm 0.71$ $93.34\pm 0.41$ $93.50\pm 0.52$ $210.79\pm 1.00$ $206.95\pm 1.34$
A-CF $54.89\pm 2.23$ $53.96\pm 1.56$ $99.48\pm 0.47$ $99.48\pm 0.32$ $93.13\pm 1.94$ $93.50\pm 1.46$ $94.09\pm 1.06$ $93.66\pm 1.56$ $85.83\pm 2.01$ $82.83\pm 1.35$
@@
Random Class Forgetting (Condensation Time for $100$ IPC: $4.90615\pm 0.1941$ s)
Retraining $51.57\pm 0.76$ $51.41\pm 0.77$ $99.97\pm 0.01$ $40.85\pm 16.59$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $85.33\pm 9.14$ $36.53\pm 16.44$ $3177.10\pm 18.08$ $71.35\pm 0.22$
CF $51.34\pm 1.03$ $46.39\pm 2.68$ $99.97\pm 0.02$ $94.73\pm 0.02$ $00.00\pm 0$ 0.00 $62.04\pm 0.09$ $85.45\pm 9.50$ $88.79\pm 0.56$ $1587.86\pm 8.52$ $35.61\pm 0.09$
Bad Distillation $48.79\pm 4.05$ $49.77\pm 8.64$ $99.91\pm 0.01$ $99.82\pm 0.38$ $0.00\pm 0.09$ $00.00\pm 0.87$ $89.99\pm 5.84$ $87.59\pm 6.84$ $1217.71\pm 6.58$ $221.71\pm 0.58$
$L_{1}$ Sparse $52.67\pm 5.92$ $59.40\pm 2.35$ $99.24\pm 0.10$ $96.64\pm 0.30$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $83.98\pm 9.43$ $83.07\pm 1.00$ $1298.78\pm 6.70$ $29.29\pm 0.06$
SCRUB $49.99\pm 0.04$ $49.97\pm 0.04$ $11.70\pm 0.70$ $11.11\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $12.08\pm 9.45$ $7.63\pm 0.13$ $679.75\pm 3.76$ $63.05\pm 0.19$
Pruning $51.56\pm 3.15$ $53.94\pm 3.66$ $99.00\pm 0.23$ $100.00\pm 0.00$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $84.31\pm 9.53$ $80.41\pm 3.55$ $647.44\pm 4.07$ $14.62\pm 0.05$
A-CF $48.98\pm 3.16$ $53.30\pm 7.61$ $97.31\pm 0.84$ $94.92\pm 0.73$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $85.92\pm 3.27$ $78.39\pm 8.36$ $211.69\pm 1.13$ $4.99\pm 0.02$
10% Uniform Forgetting (Condensation Time for 100 IPC: $0.47335\pm 0.0759$ s)
Retraining $51.68\pm 4.66$ $51.03\pm 3.50$ $99.98\pm 0.02$ $99.75\pm 0.04$ $94.62\pm 0.14$ $94.44\pm 0.14$ $95.00\pm 0.15$ $94.86\pm 0.30$ $3181.39\pm 17.57$ $3124.95\pm 19.68$
CF $54.67\pm 1.51$ $53.69\pm 2.54$ $99.97\pm 0.00$ $99.87\pm 0.03$ $97.53\pm 0.12$ $97.48\pm 0.07$ $95.22\pm 0.13$ $95.24\pm 0.11$ $1587.19\pm 9.69$ $1561.58\pm 12.57$
Bad Distillation $92.25\pm 0.98$ $92.76\pm 1.05$ $96.90\pm 0.84$ $99.71\pm 0.20$ $20.13\pm 9.61$ $22.32\pm 4.01$ $89.47\pm 1.19$ $92.16\pm 1.11$ $1218.23\pm 7.08$ $1818.03\pm 11.68$
$L_{1}$ Sparse $54.98\pm 4.00$ $54.43\pm 1.57$ $99.16\pm 0.21$ $99.50\pm 0.15$ $94.06\pm 0.97$ $95.77\pm 0.31$ $93.13\pm 1.06$ $93.82\pm 0.91$ $1299.93\pm 6.71$ $1278.57\pm 12.03$
SCRUB $53.62\pm 1.89$ $57.10\pm 2.62$ $99.17\pm 0.63$ $99.96\pm 0.04$ $93.29\pm 1.02$ $99.06\pm 0.35$ $93.58\pm 1.16$ $95.17\pm 0.31$ $681.54\pm 3.61$ $690.14\pm 5.75$
Pruning $54.42\pm 0.62$ $55.28\pm 4.82$ $99.06\pm 0.21$ $98.39\pm 1.19$ $95.03\pm 0.67$ $94.17\pm 1.87$ $93.73\pm 0.37$ $91.82\pm 3.67$ $647.93\pm 3.91$ $637.52\pm 5.97$
@@
[…]

### v3

@@ 4.2 The Efficiency and Efficacy of Blend Condensation
This dramatic speed-up does not compromise performance. As shown in Figure 2, Blend consistently yields higher test accuracy (e.g., 35.03% at IPC=10) compared to GM (33.01%), DM (32.2%), and IDM (33.1%). Furthermore, the synthetic prototypes generated by Blend are qualitatively superior. They are noticeably smoother and more class-consistent (e.g., clear airplane outlines) than the noisy, cluttered images from other methods. This demonstrates that our approach of grouping and blending feature-similar images is highly effective at producing coherent and informative exemplars.
@@
Random Class Forgetting (Condensation Time for 1000 IPC: $47.7812\pm 1.3323$ s leading to 80.0% retain dataset reduction)
Retraining $49.76\pm 2.57$ $50.51\pm 3.55$ $100.00\pm 0.00$ $75.50\pm 5.63$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $80.35\pm 2.08$ $62.47\pm 4.88$ $1141.71\pm 51.30$ $224.50\pm 0.14$
CF $0.18\pm 1.16 55$$5.39\pm 2.67$ $100.00\pm 0.00$ $96.39\pm 0.75$ $0.00\pm 0.00$ $0.30\pm 0.45$ $81.19\pm 2.40$ $77.75\pm 1.25$ $577.81\pm 55.36$ $112.32\pm 0.04$
@@
$L_{1}$ Sparse $47.32\pm 2.73$ $53.79\pm 0.88$ $99.00\pm 0.97$ $96.36\pm 0.55$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $8.42\pm 3.40 77$$7.66\pm 1.32$ $358.62\pm 28.80$ $69.88\pm 0.01$
SCRUB $50.12\pm 0.37$ $50.00\pm 0.00$ $5.28\pm 7.66 11$$1.11\pm 0.00$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $13.80\pm 6.90$ $10.00\pm 0.00$ $241.23\pm 23.22$ $61.58\pm 0.02$
Pruning $50.87\pm 3.22$ $50.80\pm 2.27$ $98.08\pm 0.73$ $100.00\pm 0.00$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $77.43\pm 2.41$ $72.65\pm 1.27$ $235.50\pm 6.20$ $46.74\pm 0.01$
A-CF $51.71\pm 2.29$ $48.25\pm 7.31$ $97.88\pm 4.39$ $98.92\pm 0.62$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $7.01\pm 3.20 77$$7.45\pm 3.63$ $75.76\pm 0.27$ $16.01\pm 0.96$
10% Uniform Forgetting (Condensation Time for 1000 IPC: $34.1780\pm 0.7403$ s leading to 39.8% retain dataset reduction)
Retraining $47\pm 0.61 50.05$$05\pm 1.64$ $100.00\pm 0.00$ $91.48\pm 0.41$ $87.48\pm 1.01$ $82.25\pm 1.01$ $86.84\pm 0.62$ $80.65\pm 0.48$ $1132.22\pm 33.71$ $664.73\pm 6.16$
@@ 4.3 Single-Round Unlearning
We benchmark single-round unlearning across utility metrics (Test/Retain Accuracy; higher is better) and privacy metrics (Forget Accuracy/MIA Score; lower is better). We analyze two orthogonal improvements: (i) our Blend dataset condensation method, which shrinks the retain set, and (ii) our loss-centric A-AMU algorithm, only applied as acceleration of $\rightarrowA-CF$ for brevity. The unlearning time in tables is separated from overhead from $k$-means partitioning and the Blend condensation (in case of “w/ DC”) for distinction.
@@
Random Class Forgetting (Condensation Time for 100 IPC: $4.90615\pm 0.1941$ s leading to 97.8% retain dataset reduction)
Retraining $50.45\pm 1.53$ $52.01\pm 5.38$ $99.99\pm 0.00$ $78.46\pm 3.28$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $82.71\pm 10.54$ $68.53\pm 8.79$ $1012.11\pm 4.18$ $22.79\pm 0.04$
CF $50.34\pm 3.97$ $51.15\pm 0.32$ $100.00\pm 0.00$ $91.37\pm 0.40$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $82.37\pm 10.30$ $81.30\pm 0.25$ $505.95\pm 2.05$ $11.36\pm 0.01$
Bad Distillation $55.38\pm 1.74$ $49.58\pm 6.42$ $8.49\pm 2.02 99$$9.69\pm 0.72$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $81.84\pm 9.47$ $83.04\pm 9.03$ $248.95\pm 1.34$ $45.15\pm 0.03$
$L_{1}$ Sparse $53.71\pm 3.01$ $52.07\pm 1.39$ $99.16\pm 0.96$ $96.97\pm 0.13$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $84.04\pm 1.87$ $94.18\pm 0.19$ $314.88\pm 3.28$ $7.14\pm 0.00$
SCRUB $50.47\pm 1.64$ $49.97\pm 0.03$ $2.93\pm 2.81 11$$1.11\pm 0.00$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $13.61\pm 8.37$ $7.37\pm 0.71$ $211.82\pm 2.00$ $19.57\pm 0.02$
Pruning $49.77\pm 2.35$ $51.90\pm 2.20$ $99.04\pm 0.18$ $100.00\pm 0.00$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $80.80\pm 10.02$ $77.36\pm 8.78$ $210.93\pm 1.00$ $4.80\pm 0.01$
A-CF $52.04\pm 2.70$ $54.09\pm 6.30$ $99.65\pm 0.59$ $97.55\pm 0.71$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $80.90\pm 9.38$ $84.49\pm 2.80$ $83.27\pm 0.48$ $2.07\pm 0.02$
10% Uniform Forgetting (Condensation Time for 100 IPC: $0.4733\pm 0.0759$ s leading to 1.6% retain dataset reduction)
Retraining $51.75\pm 1.91$ $51.61\pm 3.54$ $9.99\pm 0.00 99$$9.76\pm 0.03$ $94.60\pm 0.59$ $94.71\pm 0.26$ $95.18\pm 0.16$ $95.12\pm 0.17$ $1011.96\pm 6.45$ $992.98\pm 7.97$
@@
Bad Distillation $92.23\pm 0.43$ $93.57\pm 0.44$ $6.94\pm 1.02 99$$9.75\pm 0.11$ $23.77\pm 8.50$ $23.77\pm 2.00$ $89.73\pm 2.44$ $91.72\pm 0.41$ $248.75\pm 0.95$ $369.65\pm 2.02$
$L_{1}$ Sparse $53.55\pm 1.96$ $53.12\pm 2.44$ $9.32\pm 0.44 99$$9.47\pm 0.21$ $93.80\pm 1.00$ $94.32\pm 0.87$ $93.60\pm 0.57$ $93.33\pm 0.20$ $314.80\pm 2.90$ $308.10\pm 1.58$
SCRUB $54.99\pm 1.47$ $56.51\pm 4.66$ $66\pm 0.07 99.99$$99\pm 0.02$ $5.00\pm 0.29 99$$9.07\pm 0.44$ $94.48\pm 0.31$ $95.13\pm 0.19$ $211.64\pm 0.85$ $214.29\pm 1.76$
Pruning $54.20\pm 1.60$ $53.55\pm 0.72$ $9.33\pm 0.10 99$$9.18\pm 0.40$ $91\pm 0.16 94.94$$94\pm 0.71$ $93.34\pm 0.41$ $93.50\pm 0.52$ $210.79\pm 1.00$ $206.95\pm 1.34$
A-CF $54.89\pm 2.23$ $53.96\pm 1.56$ $9.48\pm 0.47 99$$9.48\pm 0.32$ $93.13\pm 1.94$ $93.50\pm 1.46$ $94.09\pm 1.06$ $93.66\pm 1.56$ $85.83\pm 2.01$ $82.83\pm 1.35$
@@
Random Class Forgetting (Condensation Time for 100 IPC: $4.90615\pm 0.1941$ s leading to 97.8% retain dataset reduction))
Retraining $51.57\pm 0.76$ $51.41\pm 0.77$ $99.97\pm 0.01$ $40.85\pm 16.59$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $85.33\pm 9.14$ $36.53\pm 16.44$ $3177.10\pm 18.08$ $71.35\pm 0.22$
CF $51.34\pm 1.03$ $46.39\pm 2.68$ $99.97\pm 0.02$ $94.73\pm 0.02$ $0.00\pm 0.00$ $62.04\pm 0.09$ $5.45\pm 9.50 88$$8.79\pm 0.56$ $1587.86\pm 8.52$ $35.61\pm 0.09$
Bad Distillation $48.79\pm 4.05$ $49.77\pm 8.64$ $9.91\pm 0.01 99$$9.82\pm 0.38$ $0\pm 0.09 00.00$$00\pm 0.87$ $89.99\pm 5.84$ $87.59\pm 6.84$ $1217.71\pm 6.58$ $221.71\pm 0.58$
$L_{1}$ Sparse $52.67\pm 5.92$ $59.40\pm 2.35$ $99.24\pm 0.10$ $96.64\pm 0.30$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $83.98\pm 9.43$ $83.07\pm 1.00$ $1298.78\pm 6.70$ $29.29\pm 0.06$
SCRUB $49.99\pm 0.04$ $49.97\pm 0.04$ $70\pm 0.70 11.11$$11\pm 0.00$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $12.08\pm 9.45$ $7.63\pm 0.13$ $679.75\pm 3.76$ $63.05\pm 0.19$
Pruning $51.56\pm 3.15$ $53.94\pm 3.66$ $99.00\pm 0.23$ $100.00\pm 0.00$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $84.31\pm 9.53$ $80.41\pm 3.55$ $647.44\pm 4.07$ $14.62\pm 0.05$
A-CF $48.98\pm 3.16$ $53.30\pm 7.61$ $97.31\pm 0.84$ $94.92\pm 0.73$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $85.92\pm 3.27$ $78.39\pm 8.36$ $211.69\pm 1.13$ $4.99\pm 0.02$
10% Uniform Forgetting (Condensation Time for 100 IPC: $0.47335\pm 0.0759$ s leading to 1.6% retain dataset reduction)
Retraining $51.68\pm 4.66$ $51.03\pm 3.50$ $9.98\pm 0.02 99$$9.75\pm 0.04$ $94.62\pm 0.14$ $94.44\pm 0.14$ $95.00\pm 0.15$ $94.86\pm 0.30$ $3181.39\pm 17.57$ $3124.95\pm 19.68$
CF $54.67\pm 1.51$ $53.69\pm 2.54$ $9.97\pm 0.00 99$$9.87\pm 0.03$ $97.53\pm 0.12$ $97.48\pm 0.07$ $95.22\pm 0.13$ $95.24\pm 0.11$ $1587.19\pm 9.69$ $1561.58\pm 12.57$
Bad Distillation $92.25\pm 0.98$ $92.76\pm 1.05$ $6.90\pm 0.84 99$$9.71\pm 0.20$ $0.13\pm 9.61 22$$2.32\pm 4.01$ $89.47\pm 1.19$ $92.16\pm 1.11$ $1218.23\pm 7.08$ $1818.03\pm 11.68$
$L_{1}$ Sparse $54.98\pm 4.00$ $54.43\pm 1.57$ $9.16\pm 0.21 99$$9.50\pm 0.15$ $94.06\pm 0.97$ $95.77\pm 0.31$ $93.13\pm 1.06$ $93.82\pm 0.91$ $1299.93\pm 6.71$ $1278.57\pm 12.03$
SCRUB $53.62\pm 1.89$ $57.10\pm 2.62$ $9.17\pm 0.63 99$$9.96\pm 0.04$ $3.29\pm 1.02 99$$9.06\pm 0.35$ $93.58\pm 1.16$ $95.17\pm 0.31$ $681.54\pm 3.61$ $690.14\pm 5.75$
Pruning $4.42\pm 0.62 55$$5.28\pm 4.82$ $99.06\pm 0.21$ $98.39\pm 1.19$ $95.03\pm 0.67$ $94.17\pm 1.87$ $93.73\pm 0.37$ $91.82\pm 3.67$ $647.93\pm 3.91$ $637.52\pm 5.97$
@@
Figure 4: Multi-round sequential unlearning (random 10 %) performance of without accelerated CF, L1-sparse, and SCRUB, with acceleration (A-CF, $\textbf{A-}\mathbf{L_{1}}$, A-SCRUB), with and without dataset condensation (Blend). With dataset condensation allows 34.9% retain dataset reduction